The choice magazine survey of the customers of the four biggest Australian banks found that 85% were very or fairly satisfied with their home loans. Suppose a random sample of 40 customers of the big four banks is taken. If 85% of their customers are still satisfied with their home loans, what is the probability that exactly 32 of the 40 are satisfied with their home loan? 


1
Expert's answer
2021-05-30T14:19:06-0400

By the Bernoulli formula


"\frac{40!}{32!\times8!}\times{0.85}^{32}\times{0.15}^{8}=0.109"